# Sound Field Synthesis with Acoustic Waves  [^1] Author is affiliated with Amazon Inc. USA.  [^2] This manuscript is an expanded version of a conference paper of the same title

### v1

@@ IV-C Room Impulse Response (RIR) Computation
where $\hat{\mathbf{h}}(\omega)$ is the acoustic transfer function in the frequency domain, and
@@
${\mathbf{y}}(t,\omega)=\hat{\mathbf{h}}(\omega)\ .\ u(t,\omega)$ (12)
Note that, the RIR computes only the part of the sound field that is correlated with the source signal, and it disregards the background ambient noise and other interferences in the room. Nevertheless, this ambient noise is uncorrelated with the source signal and its time-frequency map, ${\mathbf{b}}(t,\omega)$, is computed once as described in section IV-A, then added to the sound-field for any source signal. In this case, the synthesized sound-field in (12) is modified to
${\mathbf{y}}(t,\omega)=\hat{\mathbf{h}}(\omega)\ .\ u(t,\omega)+{\mathbf{b}}(t,\omega)$ (13)
IV-D Discussion
The proposed method is a combination of measurements (for room component) and simulation (for device component). A single measurement with a large microphone array is required per room position, and this measurement is reused for all devices. The measurement is processed by plane wave decomposition to compute the time-frequency map of the acoustic decomposition that is combined with device dictionary to generate the total sound field. Similarly, the device dictionary is computed once, and it is combined with any room to generate the full sound-field. The computational complexity for computing the broadband device dictionary is small because it is computed in anechoic setup. Further, it is highly parallelizable because the same process is repeated at different frequencies and at different directions for plane wave. The concept of splitting room acoustics and device acoustics significantly reduces the measurement/simulation overhead. In addition to simplifying both room and device modeling, abstracting room acoustics in a single measurement and device acoustics with the device dictionary enables reuse of either components with the other side.
@@ V Experimental Validation
The first experiment aims at validating the plane wave decomposition procedure as described in section IV-B. In Fig. 2, we show the reconstruction error of the EigenMike for two different source signals versus the number of plane wave in the expansion. The reconstruction error (aka, Goodness of Approximation, GoA) is defined as:
@@
$GoA\triangleq\frac{\int_{\omega}\|{\mathbf{y}}(\omega)-\sum_{l\in\Lambda}\alpha_{l}(\omega)\bar{\boldsymbol{\beta}}_{l}(\omega)\|^{2}}{\int_{\omega}\|{\mathbf{y}}(\omega)\|^{2}}$ (14)
where ${\mathbf{y}}(\omega)$ is the observed sound-field at the EigenMike.
Figure 2: EigenMike reconstruction error versus the number of plane waves in PWD
@@
As noted from the figure, a small number of plane waves is sufficient for sound field approximation with error less than $-20$ dB.
@@
In the following set of experiments, the RIR procedure as described in section IV-C is evaluated. The experiments were conducted in three different rooms with furniture that resemble typical bedroom and living rooms. The experiments were conducted in $24$ different positions within the three rooms. The EigenMike was colocated with four devices with different form factors and microphone array size. Two devices have cylindrical form factor, one has cube-like shape, and the fourth is a slated sphere. The $24$ test positions cover different positions in the room: middle of the room, next to a wall, and at a corner. For each position, the origin of the EigenMike and collocated device were aligned properly using laser beams. The measured and synthetic RIR are computed as described in section IV-C. In all cases, there exist strong resemblance between measured and synthetic RIR at all frequencies, and the reconstruction signal-to-noise ratio (SNR) for the devices is between $19$ to $23$ dB. An example of the transfer function and the impulse response of the measured and synthetic RIR is shown respectively in Figures 3 and 4. This is a typical behavior at all positions and devices.
@@
Figure 3: Frequency response of measured and synthetic RIR for a microphone array of size 4 on spherical cap
@@
Figure 4: Impulse response of measured and synthetic for microphone array of size 4 on spherical cap
@@
In the third set of experiments, we study the impact of mismatch between true and synthetic RIR in evaluating high level data metrics. For this test, we evaluated the False Rejection Rate (FRR) of a keyword in the source signal. The true FRR is computed by processing the device signal after the convolution of the source signal and the true RIR. Similarly, the synthetic FRR is from the convolution of the same source signal and synthetic RIR. The relative absolute FRR is defined as
Relative FRR $\text{Relative FRR Error}\triangleq\frac{|\ FRR_{true}-FRR_{synthetic}\ |}{FRR_{true}}$ (15)
@@
The cumulative density function of the relative absolute error (of all $24$ room positions and all devices) is shown in Fig. 5. From the figure, we notice that the relative error between true and synthetic FRR is less than $10\%$ with $95\%$ probability.
@@
VI Relevant Applications
VI-A Data Generation
@@
VI-B Microphone Array Processing
The traditional model for microphone array processing in the literature utilizes free-field steering vectors that capture only phase differences due to wave propagation and neglect the magnitude component due to scattering with device surface [28, 29]. This limitation results in undesirable effects with beamforming, e.g., spatial aliasing. Rather, the acoustic dictionary, as described in section III, is a generalization of free-field steering vectors that captures device response to acoustic plane wave. The entries of the acoustic dictionary provide more accurate steering vectors because they have both magnitude and phase components. Therefore, beamformer design with steering vectors from acoustic dictionary provides significantly better directivity in real room setting [3]. Further, the plane wave decomposition in (5) incorporates scattering at the device surface; hence, it provides an accurate acoustic map of the device surrounding. It can be regarded as an invertible spatial transformation that maps microphone array observations into spatial components. This enables many applications related to microphone array processing, e.g., sound source localization, sound source separation, and dereverberation.
VII Conclusions
We presented an effective procedure for sound-field synthesis based on acoustic plane waves. The system is composed of three components:
  1. 1.
    Room component: which abstracts the sound field at a point in a room as a superposition of acoustic plane waves as in (2). The complexity of room acoustics is encountered by a single recording with a large microphone array followed by plane wave decomposition to resolve the acoustic scene. Repeating the process once at multiple rooms generates a room database that is reused with any device.
  2. 2.
    Device component: which computes the response of the device surface to acoustic plane waves in the 3D space. The process is highly parallelizable and the response is computed in anechoic environment which significantly reduces the simulation complexity. The process is done once per device, and the resulting dictionary is then combined by the different rooms in the room database.
  3. 3.
    Synthesis component: which synthesizes the sound field on the device surface by combining the constituent plane waves from the first component with the surface response in the second component as in (5) .
The decoupling of room and device components significantly reduces the complexity by enabling reuse through combining the two components. We presented efficient algorithms from numerical analysis and system theory to realize the proposed system. The merits of the proposed framework and comparison with prior art were outlined in section IV-D.

### v2

@@ III-C Room Impulse Response (RIR) Computation
where $\hat{\mathbf{h}}(\omega)$ is the multichannel acoustic transfer function in the frequency domain, and
@@
$\hat{\mathbf{y}}(t,\omega)=\hat{\mathbf{h}}(\omega)\ .\ u(t,\omega)$ (12)
Note that, the RIR computes only the part of the sound field that is correlated with the source signal, and it disregards the background ambient noise and other interferences in the room. To add background and/or diffuse noise to the synthesized output, a separate time-frequency map, ${\mathbf{b}}(t,\omega)$, is computed once as described in section III-A with only background noise, then the synthesized sound-field in (12) is modified to
$\hat{\mathbf{y}}(t,\omega)=\hat{\mathbf{h}}(\omega)\ .\ u(t,\omega)+{\mathbf{b}}(t,\omega)$ (13)
III-D Discussion
The proposed method is a combination of measurements (for room component) and simulation (for device component). A single measurement with a large microphone array is required per room position, and this measurement is reused for all devices. The measurement is processed by plane wave decomposition to compute the time-frequency map of the acoustic decomposition that is combined with device dictionary to generate the total sound field. Similarly, the device dictionary is computed once, and it is combined with any room to generate the total sound-field. The computational complexity for computing the broadband device dictionary is small because it is computed in an anechoic setup. Further, it is highly parallelizable because the same process is repeated at different frequencies and at different directions for plane wave. The concept of splitting room acoustics and device acoustics significantly reduces the measurement/simulation overhead. In addition to simplifying both room and device modeling, abstracting room acoustics in a single measurement and device acoustics with the device dictionary enables reuse of either components with the other side.
@@ IV Experimental Validation
The first experiment aimed at validating the plane wave decomposition procedure as described in section III-B. In Fig. 2, we showed the reconstruction error of the EigenMike for two different source signals versus the number of plane wave in the expansion. The Goodness of Approximation, GoA, (or reconstruction error) is defined as:
@@
$\text{GoA}\triangleq\frac{\int_{\omega}\|{\mathbf{y}}(\omega)-\sum_{l\in\Lambda}\alpha_{l}(\omega)\bar{\boldsymbol{\beta}}_{l}(\omega)\|^{2}}{\int_{\omega}\|{\mathbf{y}}(\omega)\|^{2}}$ (14)
@@
where ${\mathbf{y}}(\omega)$ is the observed sound-field at the EigenMike. This was evaluated over a frequency range up to 8 kHz. As noted from the figure, a small number of plane waves is sufficient for sound field approximation with error less than $-20$ dB.
Figure 2: EigenMike reconstruction error versus the number of plane waves in the plane wave decomposition
@@
In the following set of experiments, the RIR procedure as described in section III-C was evaluated. The experiments were conducted in three different rooms with furniture that resemble typical bedrooms and living rooms, and in $24$ different positions within the three rooms. The EigenMike was placed in all positions to compute the room component that is combined with device dictionary to generate the synthetic RIR. Four other devices with different form factors and microphone array geometries were placed later at the same positions to compute the true RIR. Two devices had cylindrical form factor, one had cube-like shape, and the fourth was a slated sphere. The $24$ test positions covered different room positions: middle of the room, next to a wall, and at a corner. For each position, the origins of the EigenMike and other devices were aligned precisely using laser beams. The measured and synthetic RIR are computed as described in section III-C. In all cases, there existed strong resemblance between measured and synthetic RIR at all frequencies, and the reconstruction signal-to-noise ratio (SNR) is between $19$ and $23$ dB. An example of the transfer function and the impulse response of the measured and synthetic RIR is shown respectively in Figures 3 and 4. This is a typical behavior at all positions and devices.
@@
Figure 3: An example of frequency response of measured and synthetic RIR for a microphone array of size 4 on a slated sphere surface
@@
Figure 4: An example of impulse response of measured and synthetic for microphone array of size 4 on a slated sphere surface
@@
In the third set of experiments, we studied the impact of mismatch between true and synthetic RIR in evaluating high level data metrics. For this test, we evaluated the False Rejection Rate (FRR) of a keyword in the source signal. The true FRR was computed by processing the device signal after the convolution of the source signal and the true RIR. Similarly, the synthetic FRR was computed from the convolution of the same source signal and synthetic RIR. The relative absolute FRR is defined as
Relative FRR $\text{Relative FRR Error}\triangleq\frac{|\ \text{FRR}_{true}-\text{FRR}_{synthetic}\ |}{\text{FRR}_{true}}$ (15)
@@
The cumulative density function of the relative absolute error (of all $24$ room positions and all devices) is shown in Fig. 5. The $95$-percentile relative error between true and synthetic FRR is less than $10\%$.
@@
V Relevant Applications
V-A Data Generation
@@
V-B Microphone Array Processing
The traditional model for microphone array processing in the literature utilizes free-field steering vectors that capture only phase differences due to wave propagation and neglect the magnitude component due to scattering with device surface [29, 30]. This limitation results in undesirable effects with beamforming, e.g., spatial aliasing. Rather, the acoustic dictionary, as described in section II-B, is a generalization of free-field steering vectors that captures device response to acoustic plane wave. The entries of the acoustic dictionary provide more accurate steering vectors because they have both magnitude and phase components. Therefore, beamformer design with steering vectors from acoustic dictionary provides significantly better directivity in real room setting [3]. Further, the plane wave decomposition in (5) incorporates scattering at the device surface; hence, it provides an accurate acoustic map of the device surrounding. It can be regarded as an invertible spatial transformation that maps microphone array observations into spatial components. This enables many applications related to microphone array processing, e.g., sound source localization, sound source separation, and dereverberation.
VI Conclusions
The proposed framework enables generating accurate multichannel audio data for a general microphone array mounted on a rigid surface of an arbitrary form factor using only the CAD model of the surface and the coordinates of the microphone array. It is a balanced approach between measurement-based methods and simulation-based methods. The framework significantly reduces the cost of hardware validation and data collection. Its effectiveness is established by rigorous validation with physical data under diverse scenarios.
The key contribution of the proposed method is the decoupling of the room impact and the device-dependent impact in estimating the sound-field at the microphone array. This decoupling enables reusing the room component with different hardware designs, as if the corresponding devices are placed in the same room location. The room acoustics modeling is performed using plane wave decomposition, after collecting room measurements with a large microphone array to achieve high accuracy modeling as shown in the experimental validation. Device modeling through the device acoustic dictionary provides a general model that works with any microphone array geometry and any form factor of the surface. Though the discussion focused primarily on far-field sources and acoustic plane waves, it can be straightforwardly extended to the near field case by appending the device dictionary with acoustic spherical waves.